# JSNS2 Collaboration First Measurement of Missing Energy Due to Nuclear Effects in Monoenergetic Neutrino Charged Current Interactions

### v1

@@ II Reconstruction, Simulation, and Analysis
Figure 4 shows the observed differential cross section in terms of and for the KDAR CC interaction on carbon with statistical and systematic error bars compared to predictions from NuWro (with carbon spectral function), GiBUU, and a relativistic mean field (RMF) prediction [46, 47] using the ACHILLES event generator to simulate FSI [48, 49]. While statistical errors dominate in most of the measurement bins, uncertainties associated with the energy scale and resolution, proton Birks’ constant, and generator uncertainty (affecting the unfolding matrix) contribute significantly in the peak region. The comparisons between the data () and predictions (), using the full covariance matrix () are reported in the figure as well. For simplicity, we consider the higher-statistics region of MeV across 16 bins () when forming the . As can be seen, the generator and model predictions disagree with each other substantially and none of them is able to adequately match the data well. While it is difficult to isolate a single cause for this discrepancy across all measurement bins considered, it is much more likely that multiple uncertain nuclear effects are contributing, including FSI, the momentum and separation energy distributions of the nucleons, short-range correlations among nucleons, and even aspects of the weak interaction itself. Notably, however, the data do favor a significant FSI contribution, with a large improvement after turning on this effect within the NuWro event generator, which best matches the data overall. In general, this measurement serves as a unique and first-of-its-kind benchmark, simultaneously sensitive to multiple aspects of the event, towards elucidating low energy neutrino-nucleus interactions.
@@ III Conclusion
We have presented the first measurement of missing energy for monoenergetic CC interactions based on a % pure sample of 621 total collected KDAR candidate events. The shape-only differential cross section extracted disagrees significantly with the generator and model predictions considered. This result serves as a known-neutrino-energy, standard candle for improving our understanding of difficult to model low energy neutrino interactions at the transition between on-nucleon and on-nucleus scattering. In the future, JSNS2 and JSNS2-II [50] are poised to follow-up on this initial measurement with vastly improved statistics, better control over systematic uncertainties, in particular the proton’s Birks’ constant and energy scale/resolution, and the ability to distinguish KDAR CC events with and without neutrons.
@@ IV Acknowledgements
We deeply thank the J-PARC for their support, especially for the MLF and the accelerator groups to provide an excellent environment for this experiment. We acknowledge the support of the Ministry of Education, Culture, Sports, Science, and Technology (MEXT) and the JSPS grants-in-aid: 16H06344, 16H03967 23K13133, 24K17074 and 20H05624, Japan. This work is also supported by the National Research Foundation of Korea (NRF): 2016R1A5A1004684, 17K1A3A7A09015973, 017K1A3A7A09016426, 2019R1A2C3004955, 2016R1D1A3B02010606, 017R1A2B4011200, 2018R1D1A1B07050425, 2020K1A3A7A09080133, 020K1A3A7A09080114, 2020R1I1A3066835, 2021R1A2C1013661, NRF-2021R1C1C2003615, 2022R1A5A1030700, RS-2023-00212787 and RS-2024-00416839. Our work has also been supported by a fund from the BK21 of the NRF. The University of Michigan gratefully acknowledges the support of the Heising-Simons Foundation. This work conducted at Brookhaven National Laboratory was supported by the U.S. Department of Energy under Contract DE-AC02-98CH10886. The work of the University of Sussex is supported by the Royal Society grant no. IESnR3n170385. We also thank the Daya Bay Collaboration for providing the Gd-LS, the RENO collaboration for providing the LS and PMTs, CIEMAT for providing the splitters, Drexel University for providing the FEE circuits and Tokyo Inst. Tech for providing FADC boards.

### v2

@@ II Reconstruction, Simulation, and Analysis
The comparisons between the data () and predictions (), using the full covariance matrix () are reported in the figure as well. For simplicity, we consider the higher statistics region of =5–85 MeV across 16 bins () when forming the . As can be seen, the generator and model predictions disagree with each other substantially and none of them is able to adequately match the data well. While it is difficult to isolate a single cause for this discrepancy across all measurement bins considered, it is much more likely that multiple uncertain nuclear effects are contributing, including FSI, the momentum and separation energy distributions of the nucleons, short-range correlations among nucleons, and even aspects of the weak interaction itself. Notably, however, the data do favor a significant FSI contribution, with a large improvement after turning on this effect within the NuWro event generator, which best matches the data overall. Included as part of this publication is Supplementary Material that provides the cross-section result as well as covariance matrices for each quantified source of uncertainty. Also provided is the background subtracted measurement before unfolding, the simulation job cards used for the NuWro and GiBUU simulations, and example analysis software.
The analysis allows for events with missing energy less than zero; the observation prefers a small amount of events in that region but is consistent with zero. We also note that the primary peak occurs at MeV and the secondary peak occurs at MeV, consistent with the estimated removal energy for, respectively, p-shell and s-shell nucleons [54, 55]. In general, this measurement serves as a unique and first-of-its-kind benchmark, simultaneously sensitive to multiple aspects of the event (including what nuclear shell the neutrino interacted with), towards elucidating low energy neutrino-nucleus interactions.
@@ III Conclusion
We have presented the first measurement of missing energy for monoenergetic CC interactions based on a % pure sample of 621 total collected KDAR candidate events. The shape-only differential cross section extracted disagrees significantly with the generator and model predictions considered. This result serves as a known-neutrino-energy, standard candle for improving our understanding of difficult to model low-energy neutrino interactions at the transition between on-nucleon and on-nucleus scattering. In the future, JSNS2 and JSNS2-II [56] are poised to follow up on this initial measurement with vastly improved statistics, better control over systematic uncertainties, in particular the proton’s Birks’ constant and energy scale and resolution, and the ability to distinguish KDAR CC events with and without neutrons.
@@ IV Acknowledgements
We deeply thank the J-PARC for their support, especially for the MLF and the accelerator groups to provide an excellent environment for this experiment. We acknowledge the support of the Ministry of Education, Culture, Sports, Science, and Technology (MEXT) and the JSPS grants-in-aid: 16H06344, 16H03967 23K13133, 24K17074 and 20H05624, Japan. The work is also supported by the National Research Foundation of Korea (NRF): 2016R1A5A1004684, 17K1A3A7A09015973, 017K1A3A7A09016426, 2019R1A2C3004955, 2016R1D1A3B02010606, 017R1A2B4011200, 2018R1D1A1B07050425, 2020K1A3A7A09080133, 020K1A3A7A09080114, 2020R1I1A3066835, 2021R1A2C1013661, 2021R1A6A1A03043957, NRF-2021R1C1C2003615, 2022R1A5A1030700, RS-2023-00212787 and RS-2024-00416839. Our work has also been supported by a fund from the BK21 of the NRF. The University of Michigan gratefully acknowledges the support of the Heising-Simons Foundation. This work conducted at Brookhaven National Laboratory was supported by the U.S. Department of Energy under Contract DE-AC02-98CH10886. The work of the University of Sussex is supported by the Royal Society grant no. IESnR3n170385. We also thank the Daya Bay collaboration for providing the Gd-LS, the RENO collaboration for providing the LS and PMTs, CIEMAT for providing the splitters, Drexel University for providing the FEE circuits and Tokyo Inst. Tech for providing FADC boards. We would also like to thank Alexis Nikolakopoulos for providing the RMF+ACHILLES simulation sample to us.